OpenStudy (anonymous):

Help Please!!! The two equal angles in an isosceles triangle are each 21 degrees smaller than the third angle. What is the measure of each angle?

jimthompson5910 (jim_thompson5910):

let x = base angles (the 2 equal angles) and y = vertex angle so because "The two equal angles in an isosceles triangle are each 21 degrees smaller than the third angle", we know that x = y - 21 All angles must add to 180 degrees, so x+x+y = 180 and we can plug in x = y - 21 to get x+x+y = 180 (y-21)+(y-21)+y = 180 3y - 42 = 180 Solve for y. Then use this to find x

OpenStudy (anonymous):

Thank you very much!

jimthompson5910 (jim_thompson5910):

yw